OpenStudy (anonymous):

simplify the equation express the solution as a radical in the simplest form 4 sqrt x+3=x

OpenStudy (anonymous):

is 3 in the square root as well?

OpenStudy (anonymous):

yes it is \[4\sqrt{x+3}=x\]

OpenStudy (anonymous):

ok what is the /4 next to the problem for?

OpenStudy (anonymous):

Eh..

OpenStudy (anonymous):

You forgot to square the x.

OpenStudy (anonymous):

@angela wait---isnt it supposed to be x^2/16 since u square both sides

OpenStudy (angela210793):

:(:(

OpenStudy (anonymous):

\[\sqrt{x+3} = \frac{x}{4}\]\[\implies x+3 = \frac{x^2}{16}\qquad;\; x > -3\]

OpenStudy (anonymous):

me=??????????????? what

OpenStudy (anonymous):

\[\implies -x^2 + 16x + 48 = 0\qquad;\; x > -3\]

OpenStudy (anonymous):

can someone show me for the beginning please

OpenStudy (anonymous):

Is the problem this? \[4\sqrt{x} + 3 = x\] or this \[4\sqrt{x+3} = x\]

OpenStudy (anonymous):

You can write that as \[\sqrt{x+3}= x \div4\] . Then if you square both sides you get x+3 = x^2/16 . Do you know how to do the rest?

OpenStudy (anonymous):

@polpak yes 2nd one

OpenStudy (anonymous):

ok i got to \[x+3=x^2/16 \] now what

OpenStudy (anonymous):

multiply both sides by 16

OpenStudy (anonymous):

ok i have 16x+48=x^2

OpenStudy (anonymous):

subtract x^2 from both sides

OpenStudy (anonymous):

ok now i have -x^2+16x+48=0

OpenStudy (anonymous):

use quadratic formula.

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

Are you familiar with this ? For \[ax^2+bx+c=0\] \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

oh that ok let me try to do that

OpenStudy (anonymous):

ok yea i need help

OpenStudy (anonymous):

what is the triangle for

OpenStudy (anonymous):

in our case, what are : a=? b=? c=?

OpenStudy (anonymous):

a=x^2 b=16x c=48

OpenStudy (anonymous):

a=-x^2

OpenStudy (anonymous):

The triangle is just a symbol for the discriminant (that radical)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

a is the coefficient of x^2. in our case a=-1

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so what are: b=? c=?

OpenStudy (anonymous):

b=16x c=48

OpenStudy (anonymous):

\[x=-16x \pm \sqrt{16^2(-1)(48)}/2(-1)\]

OpenStudy (anonymous):

Note that below : a*x^2 + b*x + c = 0 a is the number in front of x^2 b is the number in front of x and c is just a number so: a=-1 b=16 c=48

OpenStudy (anonymous):

is that kinda what it is supposed to look like

OpenStudy (anonymous):

there is no "x" . a b and c are just numbers - see their values I posted above.

OpenStudy (anonymous):

oh okay i forgot to take that out but other than that was it okaay

OpenStudy (anonymous):

but let's start with the discriminant first. What is: \[\sqrt{b^2-4ac}\]

OpenStudy (anonymous):

\[\sqrt{16^2(-1)(48)}\]

OpenStudy (anonymous):

crap \[\sqrt{16^2-4(-1)(48)}\] there

OpenStudy (anonymous):

was that right?

OpenStudy (anonymous):

The discriminant is actually the value INSIDE the radical. b^2-4ac So you got b^2-4ac=16^2+4*48=448

OpenStudy (anonymous):

now what can you do with \[\sqrt{448}\] Can you take anything out of the radical ?

OpenStudy (anonymous):

im not sure what you mean

OpenStudy (anonymous):

\[\sqrt{448}=\sqrt{8^2*7}\]

OpenStudy (anonymous):

oh okay now i have done that

OpenStudy (anonymous):

so what can we take out of the radical ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

we have 8^2 inside the radical, so we can take 8 out of the radical

OpenStudy (anonymous):

ok and how

OpenStudy (anonymous):

so its 8 sqrt7 ?

OpenStudy (anonymous):

so we will end up with this: \[8\sqrt{7}\]

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

ok so now look at the whole formula again and plug in the rest of the stuff \[x=\frac{-b \pm 8\sqrt{7}}{2a}\]

OpenStudy (anonymous):

okay \[x=-16\pm8\sqrt{7}/2(-1)\] like that

OpenStudy (anonymous):

notice that the (-b) is supposed to be divided by (2a) as well

OpenStudy (anonymous):

use the draw tool :)

OpenStudy (anonymous):